trace the curve , x^2= y^2((x+1)^3 ) , stating all the points used in process


1
Expert's answer
2022-04-18T00:23:52-0400

ANSWER

Let "\Gamma" be the curve defined by the equation "x^2=y^2(x+1)^3" .

"\Gamma=\left \{ \left ( x,y \right ):x>-1, x^{2} =y^{2}\left ( x+1 \right )^{3}\right \}" ,

then "\Gamma=\Gamma_{1}\cup\Gamma_{2}" , where "\Gamma_{1}=""\left \{ \left ( x,f(x)) \right ):x>-1 \right \}" , "\Gamma_{2}= \left \{ \left ( x,-f(x)) \right ):x>-1 \right \}," "f(x)= \frac{x}{(x+1)^{\frac{3}{2}}}" . "f(0)=0"

"f'(x)=\frac { { \left( x+1 \right) }^{ \frac { 3 }{ 2 } }-\frac { 3 }{ 2 } x{ \left( x+1 \right) }^{ \frac { 1 }{ 2 } } }{ { \left( x+1 \right) }^{ 3\quad } }" "=\frac{1}{2}\cdot \frac{(2-x)}{(1+x)^\frac{5}{2}}"

"f''(x)=-\frac { 1 }{ 2 } \cdot \frac { { \quad \left( x+1 \right) }^{ \frac { 5 }{ 2 } }+\frac { 5 }{ 2 } (2-x){ \left( x+1 \right) }^{ \frac { 3 }{ 2 } } }{ { \left( x+1 \right) }^{ 5\quad } } =\frac { 3 }{ 4 } \cdot \frac { (x-4) }{ { \left( x+1 \right) }^{ \frac { 7 }{ 2 } } }"

In the interval "(-1,2)" "f'(x)>0" , so the function "f" increases. In the interval "( 2 ,+\infty)" "f'(x)<0" , so the function "f" decreases. Thus "max f(x)=f(2)=" "\frac{2}{3\sqrt{3}}" .

On the graph of the function this is point B.

"f''(x)<0" if "x\in(-1,4)" and "f''(x)>0" if "x\in(4, +\infty)" .Therefore, the function "f" in the interval "(-1,4)" is convex upwards , in the interval "(4, +\infty)" is convex downwards (see point D on the graph of the function).

"\lim _{ x\rightarrow +\infty }{ f(x)= } \lim _{ x\rightarrow +\infty }{ \frac { x }{ { x }^{ \frac { 3 }{ 2 } }{ \left( 1+\frac { 1 }{ x } \right) }^{ \frac { 3 }{ 2 } } } = } \lim _{ x\rightarrow +\infty }{ \frac { 1 }{ { x }^{ \frac { 1 }{ 2 } }{ \left( 1+\frac { 1 }{ x } \right) }^{ \frac { 3 }{ 2 } } } = } 0"

"\lim _{ x\rightarrow -1 }{ f(x)= } \lim _{ x\rightarrow -1 } \frac { x }{ { \left( 1+x \right) }^{ \frac { 3 }{ 2 } } } { =-\infty }" . Hence, "y=0" is the horizontal asymptote, "x=-1" is the vertical asymptote.